Write the equation of the circle of the smallest radius that the circle x2 +   y2    -2x-12y + 29=0 touches inside and touches line x-y-3=0 
I have no idea .. :( 

Grade:12th pass

1 Answers

Arun
5 years ago
Let centre = (x,y)
 
Now
 
r = (x-y-3 )/sqrt(2)
 
Also
 
Sqrt(1 + 36 -29) - sqrt[(x-1)² + (y-6)²] =  x-y-3/sqrt(2)
 
Solve these two equations and solve for x and y.